given the linear equation y=7/5x+2 find the y-coordinates of the points(-5,),(0,), and (5,) plot these points and graph the linear equation. please show all of your work.
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

find the x and y intercepts and graph the equation plotting the intercepts please show all of your work. 6x-7y=42
-----
x-intercept = ?
Let y = 0, then 6x = 42 and x = 7
--------------
y-intercept = ?
Let x = 0, then -7y = 42 and y = -6
=======================================
cheers,
Stan H.

Answer by mananth(16946)

y=7/5x+2 find the y-coordinates of the points(-5,),(0,), and (5,) plot these points and graph the linear equation. please show all of your work.
y=7/5x+2 x=-5, y=-5 (-5,-5)
x=0, y=2(0,2)
x=5 , y=9 (5,9)